OpenStudy (anonymous):

The physical education teacher recorded the number of sit-ups the students on the swim team were able to do at the beginning of the school year, as shown below. 11, 14, 16, 25, 14, 11, 14, 16 Which measure of central tendency best represents the students' abilities? A.) mean B.) median C.) mode

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

11, 11,14,14,14, 16, 16, 25 mean is : (11(2) + 14(3) + 16(2) + 25)/8 = 121/8 = 15.13 median is : 14 mode is : 14 I am not too sure....I would guess mean because median and mode are the same. If I were you, I would get a second opinion